OpenStudy (fernandez555):

can you convert each of these equations to slope intercept form? 1. y - 4 = ½ x - 3 2. y - 5 = 1/2(x - 5)

OpenStudy (anonymous):

1. y=1/2x+1 2. y=1/2x+2.5 (not too sure about this one)